# Disentangling Latent Shifts of In-Context Learning with Weak Supervision

### v2

@@ 4 Analysis of Weak-to-Strong Generalization
Pseudo-label correction, where the student model revises the labels predicted by the teacher, builds directly on local consistency and plays a central role in enabling W2S generalization [22]. When the model exhibits stable behavior under small input perturbations, as captured by a low Lipschitz constant, it is more likely to propagate corrections reliably across neighboring inputs in representation space. This local consistency forms the foundation for accurate correction of noisy pseudo-labels in high-confidence regions. As corrected labels accumulate, they create a foundation for generalization to nearby, low-confidence examples – gradually expanding the model’s coverage and facilitating a transition from local consistency to broader generalization. Figure 2(b) shows how the rate of corrected pseudo-labels evolves during training on GLUE datasets. As training progresses, the percentage of corrected pseudo-labels steadily increases, demonstrating wilda’s capacity to exhibit W2S generalization. Notably, the rate of pseudo-label correction plateaus faster for simpler datasets like sst and qnli, which have lower linguistic variability.
@@ 6 Conclusion
We tackled the challenges of stability and long-context handling that arise when processing multiple demonstrations in ICL within LLMs. To address these issues, we introduced wilda, a method that disentangles the latent shifts induced by demonstrations from those of the query, leveraging a teacher–student framework. wilda encodes these latent shifts into an adapter module, enabling the student model to handle queries without requiring demonstrations in the input. Moreover, wilda allows efficient handling of large demonstration sets by chunking them into manageable subsets, each processed through separate adapter modules. This not only reduces the instability caused by demonstration selection and ordering but also alleviates the context window limitations inherent in transformer-based models. We demonstrated that wilda exhibits weak-to-strong generalization by refining pseudo-labels through progressive corrections, expanding from local consistency to a more comprehensive coverage across the representation space. Our empirical evaluation of wilda showed that it consistently outperforms traditional ICL methods, significantly improving generalization and stability across diverse natural language understanding datasets.
@@ A.2 Expanding the Concatenated Matrices
A.3 Performing the Matrix Multiplication
@@ Appendix D Additional Results
D.1 Supplementary Tables
Here, we present additional results that supplement those in the main paper.
@@ D.1 Supplementary Tables
Table 6: ID generalization scores for $n$-shot scenarios ($n=4,8,32$, with $\mathcal{D}_{\text{unlab}}=100$) for Llama 3 (8B). The standard deviations of $10$ runs are shown as subscripts.
@@
D.2 Few-shot wilda
wilda is primarily designed for 0-shot operation, leveraging unlabeled data to encode task-specific information within the adapter. To examine its performance in few-shot setups, we evaluated wilda-S using Llama 3 (8B) in a $16$-shot configuration, where $16$ additional demonstrations were encoded into the adapter, resulting in a total of $32$ labeled instances. This setup was compared against standard $32$-shot ICL, as well as two wilda-S variants utilizing $32$ labeled instances in a 0-shot configuration. Additionally, we included a baseline for a 0-shot setup with only $16$ encoded demonstrations.
To standardize comparisons, we denote each wilda variant using the format $n/d$, where $n$ represents the number of shots (n-shot) and $d$ indicates the number of demonstrations encoded in the adapter. The results, averaged over 10 runs, are shown in Table 9.
Table 9: Performance comparison of wilda-S configurations and standard 32-shot ICL averaged over $10$ runs.
Method rte sst qnli mnli cola mrpc qqp math misc 32-shot ICL 75.3 93.2 77.7 69.1 58.3 76.4 74.2 43.0 84.5 wilda-S (0/32) 87.9 97.9 83.1 74.0 64.6 79.4 74.8 56.5 89.0 wilda-S (0/16) 86.0 96.1 81.4 73.1 64.3 77.7 73.1 49.5 88.0 wilda-S (16/16) 87.3 96.4 82.2 74.6 65.4 78.2 74.5 51.0 89.0
The results demonstrate that wilda-S in the $16$-shot configuration with $16$ encoded demonstrations ($16/16$) outperforms both standard $32$-shot ICL and wilda-S ($0/16$) across all datasets, showcasing its ability to utilize additional context during inference. However, it slightly underperforms compared to the 0-shot wilda-S variant with 32 encoded demonstrations ($0/32$), likely due to the fine-tuning process that is exclusive to the 0-shot setup. Nevertheless, the strong performance in n-shot setups ($n>0$) highlights the flexibility and efficacy of wilda-S in leveraging additional context provided within the prompt.
@@ E.2 Hyperparameters
  • $\textbf{r}=8$
    The rank of the low-rank matrices used to decompose the original weight matrix in LoRA. A smaller $r$ reduces the parameter count while retaining essential information.
  • $\bm{\alpha}=32$:
    A scaling factor applied to the low-rank updates, balancing the influence of the original weights and the low-rank matrices.
  • Dropout: 0.1
    The dropout rate applied to the low-rank updates.
  • Target modules:
    q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj
@@ E.3 Computing Infrastructure
Model Hugging Face ID Parameters Context window size Training tokens Adapter size Llama 3 Meta-Llama-3-8Bb 8B 8k 15T 21M Llama 2 Llama-2-7b 7B 4k 2T 20M Phi 3 Phi-3-mini-4k-instruct 3.8B 4k 3.3T 4.5M

### v3

@@ 4 Analysis of Weak-to-Strong Generalization
The student’s ability to revise the teacher’s predicted labels, known as pseudo-label correction, builds directly on local consistency and serves as a key mechanism underlying W2S generalization [23]. When the model exhibits stable behavior under small input perturbations, as captured by a low Lipschitz constant, it is more likely to propagate corrections reliably across neighboring inputs in representation space. This local consistency forms the foundation for accurate correction of noisy pseudo-labels in high-confidence regions. As corrected labels accumulate, they create a foundation for generalization to nearby, low-confidence examples – gradually expanding the model’s coverage and facilitating a transition from local consistency to broader generalization. Figure 2(b) shows how the rate of corrected pseudo-labels evolves during training on GLUE datasets. As training progresses, the percentage of corrected pseudo-labels steadily increases, demonstrating wilda’s capacity to exhibit W2S generalization. Notably, the rate of pseudo-label correction plateaus faster for simpler datasets like sst and qnli, which have lower linguistic variability.
@@
Overall, the results indicate that wilda does not merely replicate the teacher’s outputs but progressively reorganizes its representation space to reinforce locally consistent decision regions. This smoothing effect allows reliable predictions to extend beyond areas directly supported by demonstrations, consolidating weak contextual supervision into more stable and transferable task knowledge.
@@ 6 Conclusion
To tackle the challenges of stability and long-context handling that arise when processing multiple demonstrations in ICL within LLMs, we introduced wilda, a method that disentangles the latent shifts induced by demonstrations from those of the query, leveraging a teacher–student framework. wilda encodes these latent shifts into an adapter module, enabling the student model to handle queries without requiring demonstrations in the input. Moreover, wilda allows efficient handling of large demonstration sets by chunking them into manageable subsets, each processed through separate adapter modules. This not only reduces the instability caused by demonstration selection and ordering but also alleviates the context window limitations inherent in transformer-based models. Additionally, we demonstrated that wilda exhibits weak-to-strong generalization by refining pseudo-labels through progressive corrections, expanding from local consistency to a more comprehensive coverage across the representation space. Our empirical evaluation confirms these advantages, showing that wilda consistently outperforms traditional ICL methods, significantly improving generalization and stability across diverse natural language understanding datasets. Ultimately, by treating ICL as weak supervision, wilda enables parametric generalization that can flexibly incorporate contextual adaptation, turning transient in-context behaviors into persistent, reusable parameters and providing a scalable and modular foundation for stable task adaptation in LLMs.
@@ A.2 Expanding the Concatenated Matrices
A.3 Block Matrix Multiplication
@@ Appendix D Additional Results
D.1 Supplementary ID and OOD Experiments
We provide additional experimental results that complement the analyses in the main paper. Table 6 presents extended in-domain generalization results under the $16$-shot setup with $|\mathcal{D}_{\text{unlab}}|=100$, confirming that wilda consistently outperforms standard ICL and related baselines across GLUE and MMLU benchmarks. We further evaluate on the ARC-Challenge benchmark (Table 7), where wilda-S attains the strongest accuracy, demonstrating that the observed gains extend to more demanding reasoning tasks. Table 8 examines the effect of the number of demonstrations ($n\in\{4,8,32\}$), showing that wilda-S scales predictably with $n$ and remains markedly more data-efficient than standard ICL. We also assess the influence of the unlabeled query set size (Table 9), observing steady yet saturating improvements as $|\mathcal{D}_{\text{unlab}}|$ increases from $200$ to $500$. Finally, Table 10 summarizes out-of-domain generalization results, indicating that wilda preserves strong transfer and stability across heterogeneous evaluation settings.
@@ D.1 Supplementary ID and OOD Experiments
Table 7: Accuracy on the ARC-Challenge benchmark using Llama 3 (8B) with $|\mathcal{D}_{\text{unlab}}|=100$ and $n=16$ shots. The highest score is shown in bold, and the second-best score is underlined.
Method 0-shot $n$-shot PBFT ICV Batch-ICL wilda-f wilda-s wilda-r Accuracy $38.3$ $54.5$ $50.8$ $49.9$ $48.6$ $55.7$ $\mathbf{59.2}$ $56.5$
Table 8: ID generalization scores for $n$-shot scenarios ($n=4,8,32$, with $|\mathcal{D}_{\text{unlab}}|=100$) for Llama 3 (8B). The standard deviations of $10$ runs are shown as subscripts.
@@
D.2 Task Specificity and Generalization
LoRA adapters are implemented as additive residuals and do not alter the underlying model parameters. During training, we alternate between enabling the adapter (student) and disabling it (teacher), while at inference the adapter can be toggled on or off without modifying the base model. To assess whether task-specific adaptation affects general performance, we conducted additional experiments evaluating the model on unrelated, non-overlapping dataset pairs from GLUE. In each case, both $16$-shot ICL and wilda-s were exposed to demonstrations from a different domain. Results, summarized in Table 11, show that performance differences remain within one standard deviation across $10$ random seeds. This indicates that wilda-s does not degrade the model’s general zero-shot capabilities, even when an adapter trained on a specific task remains active during evaluation on a disjoint domain.
Table 11: Cross-domain impact of wilda. To test whether task-specific adaptation harms general performance, we pair unrelated GLUE datasets and evaluate whether demonstrations from one domain deteriorate performance on another. Each model variant is exposed to $16$ demonstrations from the unrelated domain. Demo denotes the dataset used for constructing demonstrations (i.e., the adapter’s training domain), while Eval indicates the unseen evaluation task. We report mean accuracy across $10$ seeds. wilda-s refers to the model with the task-specific adapter enabled at inference.
Demo Eval 0-shot 16-shot wilda-s
QNLI RTE $62.3$ $61.9_{1.2}$ $62.1_{0.9}$
SST COLA $44.6$ $4.9_{0.5} 44$$4.7_{0.4}$
MNLI QQP $61.1$ $60.9_{0.6}$ $61.0_{0.7}$
MRPC SST $79.1$ $81.4_{1.4}$ $81.3_{1.1}$
@@ E.2 Hyperparameters
  • Rank ($r=8$): Dimensionality of the low-rank matrices used to approximate the original weights, controlling parameter efficiency.
  • Scaling factor ($\alpha=32$): Multiplier that scales the low-rank updates relative to the frozen base weights.
  • Dropout: ($0.1$): Regularization applied to the low-rank updates during training.
  • Target modules: q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj.
@@ E.3 Computing Infrastructure
Model Hugging Face ID Parameters Context window size Training tokens Adapter size Llama 3 Meta-Llama-3-8B 8B 8k 15T 21M Llama 2 Llama-2-7b 7B 4k 2T 20M Phi 3 Phi-3-mini-4k-instruct 3.8B 4k 3.3T 4.5M